A steel rod hangs vertically from a fixed support. The rod is 2.0m long and 1.5mm in diameter. What mass hung from the rod will stretch it by 0.48mm ?
Answer #1


Youngs modlulus Y of a material is given Y = FL/ADL

A = area

here Y os steel is 2e11 N/m^2

DL = change of length

L = original length

F = force

so here we apply

F = YADL/L

F = mg = 2e11 * 3.14* (0.75*0.75e-6)*0.48 e-3/(2)

mg = 84.78

m = 84.78/9.8

m = 8.65 kgs
Answer #2

F = YAe/L = (20*10^10*3.14*0.75*0.75*10^-6*.00048)/(2)

m = F/g = 8.65 kg